OpenStudy (anonymous):

Mickey has a total of 36 unmatched socks red, black, brown. There are 11 unmatched red, and four times as many unmatched black as brown. How many unmatched black does he have?

OpenStudy (anonymous):

r = red b = black x = brown r + b + x = 36 r = 11 b = 4x now we sub 11 + 4x + x = 36 11 + 5x = 36 5x = 36 - 11 5x = 25 x = 5 b = 4x b = 4(5) b = 20 There are 11 unmatched red, 20 unmatched black, and 5 unmatched brown